OpenStudy (anonymous):

best answer gets a medal!!! . West deposited $15,500 in a savings account earning 2.5% interest. What was the total interest earned at the end of 8 years? A. $310.00 B. $484.38 C. $3,100.00 D. $4,843.75

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ou want to calculate the interest on $15500 at 2.5% interest per year after 8 year(s). The formula we'll use for this is the simple interest formula, or: Where: P is the principal amount, $15500.00. r is the interest rate, 2.5% per year, or in decimal form, 2.5/100=0.025. t is the time involved, 8....year(s) time periods. So, t is 8....year time periods. To find the simple interest, we multiply 15500 × 0.025 × 8 to get that: The interest is: $3100.00 Usually now, the interest is added onto the principal to figure some new amount after 8 year(s), or 15500.00 + 3100.00 = 18600.00. For example: If you borrowed the $15500.00, you would now owe $18600.00 If you loaned someone $15500.00, you would now be due $18600.00 If owned something, like a $15500.00 bond, it would be worth $18600.00 now.

OpenStudy (anonymous):

*you

OpenStudy (anonymous):

Woops, I forgot to type the formula. Here is the formula: I = P*r*t

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

so C